# Deep Reinforcement Learning Optimization for Uncertain Nonlinear Systems via Event-Triggered Robust Adaptive Dynamic Programming

### v3

@@ 1 Introduction
Learning-based methods have become a fundamental paradigm in modern engineering systems, enabling algorithms to improve performance through data-driven adaptation without relying solely on explicit mathematical models. Over the past decade, advances in machine learning—particularly in function approximation, optimization, and representation learning—have significantly expanded the capability of intelligent systems operating under uncertainty, compared to traditional analytical methods qinRealTimeRemainingUseful2023; zhangTabNetLocallyInterpretable2024; huHierarchicalTestingRabbit2025. These approaches have been increasingly adopted in control, robotics, and even generative language modelsluDeterminationSurfaceCrack2020; zhaoRealTimeObjectDetection2024; tangTemporalScaleTransformer2025; yaoLEARNINGADOPTIONUNDERSTANDING2025. However, conventional model-based techniques may be limited in their ability to handle nonlinearities, unknown disturbances, or incomplete system knowledge .
@@
Reinforcement Learning (RL) has gained attention for complex decision-making and control in uncertain, dynamic environments tangReinforcementLearningBasedApproach2024. In control engineering, RL-based methods offer a data-driven alternative to classical model-based designs. This is useful when accurate system models are difficult to obtain. Among these methods, ADP integrates RL with optimal control theory. It facilitates near-optimal control of nonlinear systems by approximating value functions and control policies through function approximators. This eliminates the need to explicitly solve the Hamilton–Jacobi–Bellman (HJB) equation lewisReinforcementLearningAdaptive2009. However, conventional ADP frameworks often rely on continuous or periodic updates to neural network parameters. These updates impose significant computational burdens and may lead to overfitting to transient disturbances or noise.
@@
Event-triggered strategies have been widely adopted in diverse control applications, including networked and embedded systems, multi-agent coordination, and resource-constrained robotic platforms onuohaStressMatrixBasedFormation2024; onuohaDiscreteTimeStressMatrixBased2024. At the meantime, the ETM has been widely employed in both control and ADP frameworks to reduce computational load (hanEventtriggeredbasedOnlineIntegral2024a; heemelsIntroductionEventtriggeredSelftriggered2012; tabuadaEventTriggeredRealTimeScheduling2007 ). Unlike time-driven schemes, ETM s update only when systems meet a state- or error-based condition. State deviation or estimation error often directly triggers updates. This approach reduces redundant updates and preserves closed-loop stability (dongEventTriggeredAdaptiveDynamic2017b; xueAdaptiveDynamicProgramming2020; onuohaStressMatrixBasedFormation2024). By limiting updates to key events, event-triggered ADP boosts efficiency and yields policies less sensitive to disturbances.
@@
Despite these advantages, engineers must ensure robustness against external disturbances and modeling uncertainties. In practice, environmental perturbations, unmodeled dynamics, nonlinear couplings, and parameter uncertainties cause disturbances. Many robust control approaches employ feedback to reduce perturbations rather than explicitly use feedforward compensation (tang2019control; tangUnmatchedDisturbanceRejection2016; tangDisturbanceObserverBasedOptimal2024). In this context, a ESO estimates the original states and accumulated interference in real time. This allows proactive compensation of parameter mismatches, unmodeled dynamics, and external perturbations in nonlinear systems s(luoAdaptiveAffineFormation2020; tangOutputTrackingUncertain2024; hanPIDActiveDisturbance2009; chenDisturbanceObserverBasedControlRelated2016; ranNewExtendedState2021c; puClassAdaptiveExtended2015; tangDisturbanceRejectionIterative2019).
inspired by the Active Disturbance Rejection Control (ADRC) philosophy (gaoScalingBandwidthparameterizationBased2003; guoConvergenceNonlinearActive2013a), provide a powerful tool:
@@
    A unified control structure incorporating ETM is developed, in which ESO-based state estimation, disturbance compensation, and controller updates occur only at triggering instants. The resulting unified composite control framework enables an aperiodic, computationally efficient implementation of output-feedback RL control. By integrating ESO-based disturbance rejection with event-triggered RL, this work establishes a unified architecture that is not addressed in the existing literature. We propose an event-triggered learning rule for a simulation-of-experience ADP scheme. In this scheme, critic and actor networks update only at triggering instants, using both instantaneous and Extrapolated Bellman Error (EBE). Unlike existing ESORL frameworks (ranReinforcementLearningBasedDisturbanceRejection2022; tangReinforcementLearningBasedOutput2024), which use continuous or periodic learning, our mechanism yields an aperiodic, data-efficient adaptation. The analysis shows practical stability and Uniform Ultimate Boundedness (UUB) are achieved without a classical PE condition (kamalapurkarConcurrentLearningbasedApproximate2016). It also avoids redundant, high-frequency updates.
@@ 2 Problem Formulation
In this paper, the nonlinear system dynamics with uncertainty is modeled as follows:
@@
Following the ADRC philosophy (hanPIDActiveDisturbance2009), the general uncertainty is transferred to a broader state:
@@
Accroding to this definition, the $n$-th subsystem dynamics can be rewritten as $\dot{x}_{n}=x_{n+1}+f_{0}(x)+g_{0}(x)u$, so that the overall system becomes an $(n+1)$th order augmented integrator chain perturbed by the unknown term $\dot{x}_{n+1}$.
@@ 3.3 Training Process
To avoid explicitly solving the HJB equation, we adopt an actor–critic architecture enhanced with an ESO-based extrapolation mechanism (ranReinforcementLearningBasedDisturbanceRejection2022). The critic network approximates the value function using a linearly parameterized structure
$V\left(\bar{x};W_{v}\right)=W_{v}^{\mathrm{T}}\phi(\bar{x}),$
@@
To enhance state-space coverage and improve robustness, an extrapolated dataset $\mathcal{X}_{E}=\left\{\xi^{i}\right\}_{i=1}^{N}$ is generated over the admissible domain, and the approximate EBE is defined as follows:
@@
where $\alpha_{c1},\alpha_{c2}>0$ are constant adaption gains and
@@ 4 Integrated Control Framework
The composite control architecture includes the uncertain. nonlinear system (1), the ESO (4), the ADP-based controller (10), and the proposed ETM. Suppose that: (i) Assumptions 1–5 hold; (ii) the observer gain vector is chosen such that the ESO error dynamics are globally asymptotically stable; (iii) the uncertainty compensation renders the plant dynamics equivalent to the nominal model used in the ADP design; and (iv) the estimation errors of the actor–critic neural-network weights are uniformly ultimately bounded. Then all closed-loop signals are uniformly ultimately bounded, and the state $x(t)$ converges to a small neighborhood of the origin.
@@ 5.1.1 Performance Analysis
As illustrated in Fig. 4, the control input produced by the composite controller stays well behaved and within bounds, even when roughly $71\%$ of its possible update instances are omitted. In contrast to the periodically updated ADP (red dashed), the ETM greatly reduces the number of control updates while maintaining comparable transient response and steady-state performance.

### v4

@@ 1 Introduction
Learning-based methods have become a fundamental paradigm in modern engineering systems, enabling algorithms to improve performance through data-driven adaptation without relying solely on explicit mathematical models. Over the past decade, advances in machine learning, particularly in function approximation, optimization, and representation learning, have significantly expanded the capability of intelligent systems operating under uncertainty, compared to traditional analytical methods (qinRealTimeRemainingUseful2023; zhangTabNetLocallyInterpretable2024; huHierarchicalTestingRabbit2025). These approaches have been increasingly adopted in control, robotics, and even generative language models (luDeterminationSurfaceCrack2020; zhaoRealTimeObjectDetection2024; tangTemporalScaleTransformer2025; yaoLEARNINGADOPTIONUNDERSTANDING2025). However, conventional model-based techniques may be limited in their ability to handle nonlinearities, unknown disturbances, or incomplete system knowledge.
@@
Reinforcement Learning (RL) has gained attention for complex decision-making and control in uncertain, dynamic environments (tangReinforcementLearningBasedApproach2024). In control engineering, RL-based methods offer a data-driven alternative to classical model-based designs. This is useful when accurate system models are difficult to obtain. Among these methods, ADP integrates RL with optimal control theory. It facilitates near-optimal control of nonlinear systems by approximating value functions and control policies through function approximators. This eliminates the need to explicitly solve the Hamilton–Jacobi–Bellman (HJB) equation (lewisReinforcementLearningAdaptive2009). However, conventional ADP frameworks often rely on continuous or periodic updates to neural network parameters. These updates impose significant computational burdens and may lead to overfitting to transient disturbances or noise.
@@
Event-triggered strategies have been widely adopted in diverse control applications, including networked and embedded systems, multi-agent coordination, and resource-constrained robotic platforms (onuohaStressMatrixBasedFormation2024; onuohaDiscreteTimeStressMatrixBased2024). Meanwhile, the ETM has been widely employed in both control and ADP frameworks to reduce computational load (hanEventtriggeredbasedOnlineIntegral2024a; heemelsIntroductionEventtriggeredSelftriggered2012; tabuadaEventTriggeredRealTimeScheduling2007). Unlike time-driven schemes, ETM s update only when systems meet a state- or error-based condition. State deviation or estimation error often directly triggers updates. This approach reduces redundant updates and preserves closed-loop stability (dongEventTriggeredAdaptiveDynamic2017b; xueAdaptiveDynamicProgramming2020; onuohaStressMatrixBasedFormation2024). By limiting updates to key events, event-triggered ADP boosts efficiency and yields policies less sensitive to disturbances.
@@
Despite these advantages, engineers must ensure robustness against external disturbances and modeling uncertainties. In practice, environmental perturbations, unmodeled dynamics, nonlinear couplings, and parameter uncertainties cause disturbances. Many robust control approaches employ feedback to reduce perturbations rather than explicitly use feedforward compensation (tang2019control; tangUnmatchedDisturbanceRejection2016; tangDisturbanceObserverBasedOptimal2024). In this context, a ESO estimates the original states and accumulated interference in real time. This allows proactive compensation of parameter mismatches, unmodeled dynamics, and external perturbations in nonlinear systems (luoAdaptiveAffineFormation2020; tangOutputTrackingUncertain2024; hanPIDActiveDisturbance2009; chenDisturbanceObserverBasedControlRelated2016; ranNewExtendedState2021c; puClassAdaptiveExtended2015; tangDisturbanceRejectionIterative2019).
@@
    A unified control structure incorporating ETM is developed, in which ESO-based state estimation, disturbance compensation, and controller updates occur only at triggering instants. The resulting unified composite control framework enables an aperiodic and computationally efficient realization of output-feedback RL control. By integrating ESO-based disturbance rejection with event-triggered RL, this work establishes a unified architecture not addressed in existing literature.
  2. 2.
    We propose an event-triggered learning rule for a simulation-of-experience ADP scheme, where critic and actor networks are updated solely at triggering instants using both instantaneous and Extrapolated Bellman Error (EBE). In contrast to existing ESORL frameworks (ranReinforcementLearningBasedDisturbanceRejection2022; tangReinforcementLearningBasedOutput2024), which employ continuous or periodic learning, the proposed mechanism produces an aperiodic, data-efficient adaptation. The associated analysis shows that practical stability and Uniform Ultimate Boundedness (UUB) are achieved without imposing a classical PE condition (kamalapurkarConcurrentLearningbasedApproximate2016), while avoiding redundant high-frequency learning updates.
@@ 2 Problem Formulation
In this paper, the nonlinear system dynamics with uncertainty are modeled as follows:
@@
Following the Active Disturbance Rejection Control (ADRC) philosophy (hanPIDActiveDisturbance2009), the general uncertainty is transferred to a broader state:
@@
According to this definition, the $n$-th subsystem dynamics can be rewritten as $\dot{x}_{n}=x_{n+1}+f_{0}(x)+g_{0}(x)u$, so that the overall system becomes an $(n+1)$th order augmented integrator chain perturbed by the unknown term $\dot{x}_{n+1}$.
@@ 3.3 Training Process
To avoid explicitly solving the HJB equation, we adopt an actor–critic architecture enhanced with an ESO-based extrapolation mechanism (ranReinforcementLearningBasedDisturbanceRejection2022). The critic network approximates the value function using a linearly parameterized structure.
$V\left(\bar{x};W_{c}\right)=W_{c}^{\mathrm{T}}\phi(\bar{x}),$
@@
To enhance state-space coverage and improve robustness, an extrapolated dataset $\mathcal{X}_{E}=\left\{\xi^{i}\right\}_{i=1}^{N}$ is generated in the admissible domain, and the approximate EBE is defined as follows:
@@
where $\alpha_{c1},\alpha_{c2}>0$ are constant adaptation gains and
@@ 4 Integrated Control Framework
The composite control architecture includes the uncertain nonlinear system (1), the ESO (4), the ADP-based controller (10), and the proposed ETM. Suppose that: (i) Assumptions 1–4 hold; (ii) the observer gain vector is chosen such that the ESO error dynamics are globally asymptotically stable; (iii) the uncertainty compensation renders the plant dynamics equivalent to the nominal model used in the ADP design; and (iv) the estimation errors of the actor–critic neural-network weights are uniformly ultimately bounded. Then all closed-loop signals are uniformly ultimately bounded, and the state $x(t)$ converges to a small neighborhood of the origin.
@@ 5.1.1 Performance Analysis
As illustrated in Fig. 4, the control input produced by the composite controller stays well behaved and within bounds, even when roughly $71\%$ of its possible update instances are omitted. In contrast to the periodically updated ADP (red dashed), ETM greatly reduces the number of control updates while maintaining comparable transient response and steady-state performance.